# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v1

@@ I-A Related Works
As of the aggregation strategies in related literature, by leveraging median [16], trimmed mean [19], or iterative filtering [20], the associated FL system can tolerate a small fraction of Byzantine users. To be more robust to Byzantine attacks, some methods such as Krum [21], Federated Learning Over the Air (FLOA) [22], Robust Stochastic Aggregation (RSA) [23], and Robust Aggregating Normalized Gradient Method (RANGE) [24], have been proposed. Specifically, Krum selects a stochastic gradient as the global one, which has the shortest cumulative distance from a group of stochastic gradients that are most closely distributed, FLOA requires each user to normalize the associated local gradient to be a vector with zero mean and unit variance before offloading it to the central server, RSA penalizes the difference between local model parameters and global model parameters to isolate Byzantine users, and RANGE takes the coordinate-wise median of uploaded stochastic gradients as the aggregated one.
@@ II Problem Statement
[…]
@@ Appendix A Proof of Theorem 1
Summarizing the inequality in (A) for $t=1,2,...,T$ and dividing the summarized inequality with $\sum_{t=1}^{T}{\eta^{t}}/{2}$ for both sides, we obtain
@@ Appendix B Proof of Theorem 2
Applying the inequality in (B) for $t=1,2,...,T$, and recalling that $\gamma^{t}=\frac{1-2\mu\eta^{t}+L^{2}(\eta^{t})^{2}}{1-\lambda^{t}}$ and the definition of $q^{t}$, we then have
@@
With the help of Lemma 1, and combining (B) and (35), there is
@@ Appendix C Proof of Lemma 1
where the inequality (39c) holds because of the Cauchy-Schwarz inequality $(a_{1}+a_{2}+a_{3}+\dots+a_{n})^{2}\leqslant n(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+\dots+a_{n}^{2})$. With Assumptions 3 and 4, there are $\mathbb{E}\left\{\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\}=0$ and $\mathbb{E}\left\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\rVert^{2}\right\}\leqslant\sigma^{2}$, then (C) leads to
@@
To further upper bound the right-hand side of (C), since $\theta_{m}^{t,k}=\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2}$, we only need to inspect the bound of $\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$, which is given as follows
@@
Combining (20), (C), (C) and (C), we obtain the following inequality

### v2

@@ I-A Related Works
As of the aggregation strategies in related literature, by leveraging median [16], trimmed mean [19], or iterative filtering [20], the associated FL system can tolerate a small fraction of Byzantine users. To be more robust to Byzantine attacks, some methods such as Krum [21], Federated Learning Over the Air (FLOA) [22], Robust Stochastic Aggregation (RSA) [23], and Robust Aggregating Normalized Gradient Method (RANGE) [24], have been proposed. Specifically, Krum selects a stochastic gradient as the global one, which has the shortest cumulative distance from a group of stochastic gradients that are most closely distributed, FLOA requires each user to normalize the associated local gradient to be a vector with zero mean and unit variance before offloading it to the central server, RSA penalizes the difference between local model parameters and global model parameters to isolate Byzantine users, and RANGE takes the coordinate-wise median of uploaded stochastic gradients as the aggregated one. Other strategies try to filter out the messages from Byzantine attackers before aggregation at the central server through various measures, e.g., [28] leverges semi verified mean estimation and [29] utilizes a multi-dimensional Huber loss function.
@@ II Problem Statement
$\displaystyle\text{Var}(\nabla F_{m}(\bm{w};\xi_{m}))\leqslant\sigma^{2},\forall m\in\mathcal{M}.$ (9)
@@ Appendix A Proof of Theorem 1
Summarizing the inequality in (31) for $t=1,2,...,T$ and dividing the summarized inequality with $\sum_{t=1}^{T}{\eta^{t}}/{2}$ for both sides, we obtain
@@ Appendix B Proof of Theorem 2
Applying the inequality in (34) for $t=1,2,...,T$, and recalling that $\gamma^{t}=\frac{1-2\mu\eta^{t}+L^{2}(\eta^{t})^{2}}{1-\lambda^{t}}$ and the definition of $q^{t}$, we then have
@@
With the help of Lemma 1, and combining (33) and (35), there is
@@ Appendix C Proof of Lemma 1
where the inequality (39c) holds because of the Cauchy-Schwarz inequality $(a_{1}+a_{2}+a_{3}+\dots+a_{n})^{2}\leqslant n(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+\dots+a_{n}^{2})$. With Assumptions 3 and 4, there are $\mathbb{E}\left\{\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\}=0$ and $\mathbb{E}\left\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\rVert^{2}\right\}\leqslant\sigma^{2}$, then (39) leads to
@@
To further upper bound the right-hand side of (40), since $\theta_{m}^{t,k}=\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2}$, we only need to inspect the bound of $\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$, which is given as follows
@@
Combining (20), (C), (40) and (41), we obtain the following inequality